# Anti-rotors and algebra invariants

### v37

@@ 8.2. Type II anti-rotors
From Proposition 8.6, we know that $s^{5}$ is uncurled by the third-order anti-rotor members. But now let us assume that all orders of Type II anti-rotors are identical. In that case, $\mu^{\prime}$ also uncurls $s^{5}$. From the commutativity of the exterior derivative and a pullback, it would then follow from (8.7) that $\mu^{\prime}$ uncurls $s^{2}$. We can repeat this type of argument for all $s^{j}$ with $j$ a positive even number. Thus, equivalence of Type II anti-rotors of all orders would imply that the unique Type II anti-rotor uncurls all vector fields $s^{j}$, $j\in\mathbb{N}$, $j>1$.
To show that all orders of Type II anti-rotors are identical, according to Definition 8.5, it is “only” necessary to show that all $\mathcal{D}_{k}$ are symmetrized by being left-multiplied by the same parametrized matrix. Unfortunately, we don’t have a proof of that assertion either, so the computation evidence cited below will only leave us with,
@@
Clearly, our program would be greatly simplified if there is a unique Type I anti-rotor, and there is a unique Type II anti-rotor, and these are equal to each other. Although those possibilities might seem fanciful, consider the following sample of available computational evidence.
As with the Type I anti-rotors, we can use the algebras in Table III to examine Type II anti-rotors. In particular, it is easy to show that for every associative algebra in Table III, its second-order Type II anti-rotor is identical to its Type I anti-rotor associated with $f(s)=s^{-1}$ - as well being identical to the Type I anti-rotors associated with $f(s)=s^{j}$, for small $|j|$, $j\in\mathbb{Z}$, $j\neq 0,1$, e.g., with $f(s)=s^{3}$. For example, we can illustrate this with the noncommutative algebra in the twelfth row of Table III (being noncommutative, we have, $\hat{s}_{L}\neq\hat{s}_{R}$). For this algebra, we obtain,
@@
where $M_{\mathfrak{u}}$ is the Type I anti-rotor associated with $f(s)=s^{-1}$ (i.e., the anti-rotor as forwarded by Definition 3.1), which is the same as the Type I anti-rotor associated with $f(s)=s^{3}$. Equation (8.3) is then easily verified using the parametrized $M_{\mathfrak{u}}$ as a substitute for $\mu^{\prime}$ - which means that the second-order Type II anti-rotor is identical to $M_{\mathfrak{u}}$, since the second-order Type II anti-rotor is a subset of a corresponding Type I anti-rotor, and in this example we have shown that every member of the Type I anti-rotor associated with $f(s)=s^{-1}$ (or $f(s)=s^{3}$) is also a member of the second-order anti-rotor (i.e., it satisfies (8.3)). Without much effort, using the more general (8.6) with $k=3$, it is seen that this also holds for the third-order anti-rotor, i.e., the third-order anti-rotor is equal to $M_{\mathfrak{u}}$.

### v38

@@ 8.2. Type II anti-rotors
From Proposition 8.6, we know that $s^{5}$ is uncurled by the third-order anti-rotor members. But now let us assume that all orders of Type II anti-rotors are identical. In that case, $\mu^{\prime}$ also uncurls $s^{5}$. From the commutativity of the exterior derivative and a pullback, it would then follow from (8.7) that $\mu^{\prime}$ uncurls $s^{2}$. We can repeat this type of argument for all $s^{j}$ with $j$ a positive even number. Thus, we have
Corollary 8.7.
For a given algebra, if its Type II anti-rotors of all orders are identical then this unique Type II anti-rotor uncurls all vector fields $s^{j}$, $j\in\mathbb{N}$.
To show that all orders of Type II anti-rotors are identical, according to Definition 8.5, it is “only” necessary to show that all $\mathcal{D}_{k}$ are symmetrized by being left-multiplied by the same parametrized matrix. Unfortunately, we don’t have a proof of that assertion either, so the computational evidence cited below will only leave us with,
@@
Clearly, our program would be greatly simplified if there is a unique Type I anti-rotor, and there is a unique Type II anti-rotor, and these are equal to each other. Although those possibilities might seem fanciful, consider the following sample of easily generated computational evidence.
As with the Type I anti-rotors, we can use the algebras in Table III to examine Type II anti-rotors. In particular, it is easy to show that for every associative algebra in Table III, its second-order Type II anti-rotor is identical to its Type I anti-rotor associated with $f(s)=s^{-1}$ - which as noted earlier is identical to the Type I anti-rotors associated with $f(s)=s^{j}$, for small $|j|$, $j\in\mathbb{Z}$, $j\neq 0,1$, e.g., with $f(s)=s^{3}$. For example, we can illustrate this with the noncommutative algebra in the twelfth row of Table III (being noncommutative, we have, $\hat{s}_{L}\neq\hat{s}_{R}$). For this algebra, we obtain,
@@
where $M_{\mathfrak{u}}$ is the Type I anti-rotor associated with $f(s)=s^{-1}$ (i.e., the anti-rotor as forwarded by Definition 3.1), which (in this and the other examples from Table III) is the same as the Type I anti-rotor associated with $f(s)=s^{3}$. Equation (8.3) is then easily verified using the parametrized $M_{\mathfrak{u}}$ as a substitute for $\mu^{\prime}$ - which means that the second-order Type II anti-rotor is identical to $M_{\mathfrak{u}}$, since the second-order Type II anti-rotor is a subset of a corresponding Type I anti-rotor, and in this example we have shown that every member of the Type I anti-rotor associated with $f(s)=s^{-1}$ (or $f(s)=s^{3}$) is also a member of the second-order anti-rotor (i.e., it satisfies (8.3)). Without much effort, using the more general (8.6) with $k=3$, it is seen that this also holds for the third-order Type II anti-rotor, i.e., the third-order Type II anti-rotor is also equal to $M_{\mathfrak{u}}$.